OpenStudy (anonymous):

How many moles of copper are in 6,000,000 atoms of copper? I know I need to use the paranthesis and I know to divide by 6.02x1023. But the the answer that is listed correct is 9.96 x10^-19. I got exponent 17.

sam (.sam.):

\[6,000,000 \cancel{atoms} \times \frac{1mole}{6.02 \times 10^{23}\cancel{atoms}}=?\]